# Composite Indicator-Guided Infilling Sampling for Expensive Multi-Objective Optimization

### v1

@@ IV-A Experimental Settings
    Benchmarks Problems: Benchmarks Problems: The DTLZ [40], ZDT [41], and MaF [42] test suites are selected as the benchmark functions for our research. Referring to [38] for EMOPs, the number of decision variables ($d$) is 8 for the two-objective DTLZ and ZDT test suites and 6 for the three-objective DTLZ test suite. Referring to [32] for EMaOPs, algorithms are tested on 3, 5, 10-objective MaF problems, where the number of decision variables is 10 for all test problems except for the 10-objective MaF11 problem whose dimension is set to 11.
@@
    Performance Evaluation of Algorithms: We adopted inverted generational distance plus (IGD+) [43] and hypervolume (HV) [44] as performance indicators to evaluate the performance of algorithms. Wilcoxon’s rank-sum test [45] is conducted at a 5% significance level to assess whether the proposed CI-EMO exhibits statistically significant differences compared to the algorithms under evaluation. Specifically, the symbols “--”, “$+$”, and $\approx$” indicate the compared algorithm performs significantly worse, better, or similarly to CI-EMO, respectively.
TABLE I: IGD+ STATISTIC RESULTS (MEAN AND STANDARD DEVIATION) FOR CI-EMO AND OTHER FIVE SAEAS ALGORITHMS ON BENCHMARK PROBLEMS
@@
IV-B Comparison with state-of-the-art SAEAs
@@
To evaluate the effectiveness of CI-EMO, we compared CI-EMO against five state-of-the-art expensive multi-objective optimization algorithms. The experimental results are summarized in Table I and S-I. Table I reports the mean IGD+ values and standard deviations (std.) of the approximation Pareto-optimal solutions obtained by the six algorithms. From these results, it is clear that CI-EMO achieves the best mean IGD+ values on two-objective problems such as DTLZ1-2, DTLZ4-5, DTLZ7, ZDT1-3, as well as three-objective problems including DTLZ4 and DTLZ7. CI-EMO outperforms the compared algorithms on the majority of test problems, and CI-EMO significantly outperforms K-RVEA on 18 out of 19 problems expected ZDT4. CI-EMO performs better than K-RVEA, KTA2, EMMOEA, DirHV-EGO, and R2/D-EGO on 18, 12, 9, 15, and 14 out of 19 problems, respectively, while being inferior to these algorithms on only 0, 4, 4, 3, and 3 problems, respectively. Table S-I shows the mean HV values and standard deviations of the approximation Pareto-optimal solutions obtained by the six algorithms. CI-EMO performs better than K-RVEA, KTA2, EMMOEA, DirHV-EGO, and R2/D-EGO on 13, 7, 7, 10, and 10 out of 19 problems, while it is inferior to them only on 0, 1, 0, 3, and 3 problems, respectively. Its consistent superiority in terms of both IGD+ and HV metrics across a range of benchmark problems demonstrates its effectiveness and robustness. Thus, we can conclude that CI-EMO demonstrates strong competitiveness in solving expensive multi-objective optimization problems.
@@
To gain further insights, we analyzed the performance of CI-EMO on problems with varying Pareto front characteristics. CI-EMO performs effectively on convex Pareto fronts (e.g., ZDT1, DTLZ2), concave Pareto fronts (e.g., ZDT2), and discontinuous Pareto fronts (e.g., ZDT3, DTLZ7). As illustrated in Fig. 3, CI-EMO demonstrates superior performance in identifying well-distributed Pareto solution sets compared to other advanced methods, particularly for problems with discontinuous Pareto fronts, such as DTLZ7 and ZDT3. However, we also observe that complex problems with multimodal functional landscapes, such as ZDT4 and DTLZ3, pose significant challenges in finding the Pareto Front within a limited number of fitness evaluation budgets. IGD+ value of all methods are over 20 on ZDT4. The non-dominated solution set is still far from the Pareto frontier. The advantage of CI-EMO sampling that takes distribution into account is not exploited, resulting in CI-EMO achieving second place on this problem rather than the best result. To demonstrate the performance of CI-EMO more convincingly, we plot the convergence curves of the IGD+ values of the six algorithms on the 2-objective DTLZ, ZDT test set in Fig. 4, which are averaged over 21 independent runs. The legend is shown in the DTZL1 subfigure. From Fig. 4, we can see that when the number of function evaluations is less than 160, the proposed CI-EMO method converges relatively converges much faster for many problems, such as DTLZ1, DTLZ3, DTLZ5, DTLZ7, and ZDT1-4.
@@
TABLE II: IGD+ STATISTIC RESULTS (MEAN AND STANDARD DEVIATION) FOR CI-EMO AND OTHER FIVE SAEAS ALGORITHMS ON MaF PROBLEMS
@@
IV-C Comparison with state-of-the-art SAEAs on Many-objective Problems
@@
To further research CI-EMO performance on expensive many-objective problems, we compare algorithms on 3-, 5-, and 10-objective MaF test problems. IGD+ Results are shown in Table II. CI-EMO performs better than K-RVEA, KTA2, EMMOEA, DirHV-EGO, and R2/D-EGO on 29, 15, 15, 26, and 25 out of 33 problems, respectively, while being inferior to these algorithms on only 1, 5, 10, 4, and 4 problems, respectively. We find that CI-EMO is most efficient in solving MaF1-3 and MaF7-11. CI-EMO is weaker than EMMOEA on MaF4, MaF6, and MaF12-13 since EMMOEA has an advantage in solving concave PF problems, which has been described on [32]. Overall, CI-EMO has the best performance, which verifies the effectiveness of composite indicator-based sampling in high-dimensional objective space. Thus, experiments demonstrate CI-EMO not only has strong competitiveness in solving expensive multi-objective optimization problems but also works well on many-objective optimization problems.
IV-D Effectiveness of Composite Indicator
@@
To validate the effectiveness of the composite indicator, this paper designs two types of experiments: one to analyze the properties of each component in the composite indicator, and another to assess the impact of different components of the indicator by ablation experiments. First, CI-EMO is compared with four variants, where $rand$-EMO refers to the method of randomly selecting an individual from candidate population for real fitness evaluation, while the other three variants correspond to using one of the three performance indicators for candidate selection, respectively. “M” represents the number of objectives and "D" means the dimension of the problem. $\mathbf{I}_{1}$-EMO means only $\mathbf{I}_{1}$ is used in the composite indicator. The experimental results are presented in Table III. As shown in the table, CI-EMO significantly outperforms $rand$-EMO on 13 test problems. Furthermore, the algorithm’s performance is inferior to CI-EMO when only a single indicator ($\mathbf{I}_{1}$, $\mathbf{I}_{2}$, or $\mathbf{I}_{3}$) is used. It is evident that using different indicators produces significantly varied results across different test problems, as shown in Fig. S-1. Compared with single sampling indicator-based variants, CI-EMO has better robustness. Among the three variants that use a single performance indicator, $\mathbf{I}_{1}$-EMO performs the best. It shows no significant difference compared to CI-EMO in 14 test cases and is significantly worse in five test problems. $\mathbf{I}_{2}$-EMO, performs significantly worse than CI-EMO in 9 test cases. It performs well on the ZDT test set but underperforms on the DTLZ test set. $\mathbf{I}_{3}$-EMO exhibits the poorest performance; except for DTLZ1 and DTLZ3, it is significantly worse than CI-EMO in most of the problems. As shown in Fig. S-1, if only $\mathbf{I}_{3}$ is used, solutions on PF will lack diversity, selection of solutions based on the convergence indicator is biased towards certain regions of the PF as results on DTLZ2 and ZDT3. In addition, the algorithm benefits from multiple iterations of surrogate model-assisted NSGA-III during the generation process of the candidate population. Individuals in the candidate population have exhibited convergence. Therefore, $\mathbf{I}_{1}$-EMO and $\mathbf{I}_{2}$-EMO have better performance than $\mathbf{I}_{3}$-EMO.
@@
On the other hand, ablation experiments have been conducted. We delete one metric from the composite indicator in each variant. Results have been shown in Table IV. CI-EMO-no-$\mathbf{I}_{1}$ means that $\mathbf{I}_{2}$ and $\mathbf{I}_{3}$ are used in the composite indicator, and $\mathbf{I}_{1}$ is deleted from the composite indicator. We can see that no matter which of the three metrics is deleted, the performance of the algorithm will drop significantly. Among them, deleting $\mathbf{I}_{3}$ has the least impact on the proposed algorithm, and the effect of CI-EMO-no-$\mathbf{I}_{3}$ is better on some problems, such as ZDT1-3.
@@
TABLE III: IGD+ STATISTIC RESULTS (MEAN AND STANDARD DEVIATION) FOR CI-EMO AND OTHER FOUR VARIANTS ON BENCHMARK PROBLEMS
@@
TABLE IV: IGD+ STATISTIC RESULTS (MEAN AND STANDARD DEVIATION) FOR CI-EMO AND OTHER THREE VARIANTS ON BENCHMARK PROBLEMS
@@
TABLE V: IGD+ STATISTIC RESULTS (MEAN AND STANDARD DEVIATION) FOR CI-EMO AND THE VARIANT WITHOUT NORMALIZATION
@@
IV-E Effectiveness of normalization
For integrating the three kinds of performance metrics of convergence, diversity, and distribution, the algorithm normalizes different indicators and then superimposes them together. To verify the impact of normalization on each indicator, we used a new variant CI-EMO-no-Norm as a control. In CI-EMO-no-Norm, $\mathbf{I}_{1}$, $\mathbf{I}_{2}$, and $\mathbf{I}_{3}$ are obtained without normalization, that is, they are equal to $\theta(\boldsymbol{x})$, $d_{c}\left(\boldsymbol{x}\right)$, and $-d_{z}\left(\boldsymbol{x}\right)$, respectively. The experimental result is shown in Table V. CI-EMO significantly outperformed CI-EMO-no-Norm on 11 test questions, and there was no significant difference on the remaining 8 test questions. It shows that normalization can help combine different scale metrics and achieve the balance of convergence, diversity, and distribution.
TABLE VI: IGD+ STATISTIC RESULTS (MEAN AND STANDARD DEVIATION) FOR CI-EMO AND OTHER FOUR VARIANTS WITH DIFFERENT SAMPLING NUMBER
@@ IV-E Effectiveness of normalization
IV-F Effect of random weights of indicators
A fixed weight setting may cause the algorithm’s sampling points to exhibit a single preference and lack diversity. To address this, random weights are assigned to the three sampling indicators during their integration. To analyze the impact of random weights, we compare CI-EMO with its variant, CI-EMO-SW, where all weights are fixed and set to 1. The results are presented in Table S-II. Experimental comparisons reveal that the overall statistical results of fixed weights and random weights are not significantly different. However, random weights yield more optimal results, as highlighted in Table S-II. Furthermore, as shown in Fig. 5, for the discontinuous Pareto front problem ZDT3, using random weights demonstrates better robustness in identifying each segment of the discontinuous Pareto front. Experiments show random weights enhance, improve the diversity of sampling, and prevent issues due to fixed weights.
@@ IV-F Effect of random weights of indicators
IV-G Effect of sampling number setting each generation
In expensive optimization, the sampling number in each iteration is commonly an important parameter for algorithm performance. Some previous methods use multiple sampling strategies to select multiple candidate solutions from the candidate population at each iteration for real fitness evaluation, and each strategy is used for different purposes [29, 31]. In CI-EMO, we use a composite indicator to select only one candidate in each iteration, where the selected candidate balances the performance of convergence, diversity, and distribution.
To verify the impact of the sampling number of CI-EMO in this paper, we compared CI-EMO with four variants, where CI-EMO-q10 means the algorithm selects 10 candidate solutions for real fitness evaluation in each iteration. Similarly, CI-EMO-q5, CI-EMO-q3, and CI-EMO-q2 respectively indicate that 5, 3, and 2 candidate solutions are selected from the candidate population for real fitness value evaluation in each iteration. When the former candidate solution is selected based on the composite indicator, the candidate solution is deleted from the candidate population, and selects the next new candidate solution. As shown in Table VI, CI-EMO has the best performance. The performance of CI-EMO-q2 and CI-EMO-q3 are similar to CI-EMO. CI-EMO-q3 performs significantly worse than CI-EMO on only 3 test problems. As the number of samples continues to increase, the performance of the algorithm decreases significantly. This is because some unnecessary candidate solutions in the population are selected, making the use of the true fitness value evaluation times no longer efficient.
IV-H Computational Complexity Analysis

### v2

@@ 4.1 Experimental Settings
    Benchmarks Problems: The DTLZ (Deb et al., 2002), ZDT (Deb, 1999), and MaF (Cheng et al., 2017) test suites are selected as the benchmark functions for our research. Referring to (Zhao and Zhang, 2023) for EMOPs, the number of decision variables ($d$) is 8 for the two-objective DTLZ and ZDT test suites and 6 for the three-objective DTLZ test suite. Referring to (Qin et al., 2023) for EMaOPs, algorithms are tested on 3, 5, 10-objective MaF problems, where the number of decision variables is 10 for all test problems except for the 10-objective MaF11 problem whose dimension is set to 11.
@@
    Performance Evaluation of Algorithms: We adopted inverted generational distance plus (IGD+) (Ishibuchi et al., 2015) and hypervolume (HV) (Zitzler and Thiele, 1999) as performance indicators to evaluate the performance of algorithms. Wilcoxon’s rank-sum test (Wilcoxon et al., 1970) is conducted at a 5% significance level to assess whether the proposed CI-EMO exhibits statistically significant differences compared to the algorithms under evaluation. Specifically, the symbols “--”, “$+$”, and $\approx$” indicate the compared algorithm performs significantly worse, better, or similarly to CI-EMO, respectively.
Table 1: IGD+ STATISTIC RESULTS (MEAN AND STANDARD DEVIATION) FOR CI-EMO AND OTHER FIVE SAEAS ALGORITHMS ON BENCHMARK PROBLEMS
@@
4.2 Comparison with state-of-the-art SAEAs
@@
To evaluate the effectiveness of CI-EMO, we compared CI-EMO against five state-of-the-art expensive multi-objective optimization algorithms. The experimental results are summarized in Table 1 and S-I. Table 1 reports the mean IGD+ values and standard deviations (std.) of the approximation Pareto-optimal solutions obtained by the six algorithms. From these results, it is clear that CI-EMO achieves the best mean IGD+ values on two-objective problems such as DTLZ1-2, DTLZ4-5, DTLZ7, ZDT1-3, as well as three-objective problems including DTLZ4 and DTLZ7. CI-EMO outperforms the compared algorithms on the majority of test problems, and CI-EMO significantly outperforms K-RVEA on 18 out of 19 problems expected ZDT4. CI-EMO performs better than K-RVEA, KTA2, EMMOEA, DirHV-EGO, and R2/D-EGO on 18, 12, 9, 15, and 14 out of 19 problems, respectively, while being inferior to these algorithms on only 0, 4, 4, 3, and 3 problems, respectively. Table S-I shows the mean HV values and standard deviations of the approximation Pareto-optimal solutions obtained by the six algorithms. CI-EMO performs better than K-RVEA, KTA2, EMMOEA, DirHV-EGO, and R2/D-EGO on 13, 7, 7, 10, and 10 out of 19 problems, while it is inferior to them only on 0, 1, 0, 3, and 3 problems, respectively. Its consistent superiority in terms of both IGD+ and HV metrics across a range of benchmark problems demonstrates its effectiveness and robustness. Thus, we can conclude that CI-EMO demonstrates strong competitiveness in solving expensive multi-objective optimization problems.
@@
To gain further insights, we analyzed the performance of CI-EMO on problems with varying Pareto front characteristics. CI-EMO performs effectively on convex Pareto fronts (e.g., ZDT1, DTLZ2), concave Pareto fronts (e.g., ZDT2), and discontinuous Pareto fronts (e.g., ZDT3, DTLZ7). As illustrated in Fig. 3, CI-EMO demonstrates superior performance in identifying well-distributed Pareto solution sets compared to other advanced methods, particularly for problems with discontinuous Pareto fronts, such as DTLZ7 and ZDT3. However, we also observe that complex problems with multimodal functional landscapes, such as ZDT4 and DTLZ3, pose significant challenges in finding the Pareto Front within a limited number of fitness evaluation budgets. IGD+ values are over 20 on ZDT4. The non-dominated solution set is still far from the Pareto frontier. The advantage of CI-EMO sampling that takes distribution into account is not exploited, resulting in CI-EMO achieving second place on this problem rather than the best result. To demonstrate the performance of CI-EMO more convincingly, we plot the convergence curves of the IGD+ values of the six algorithms on the 2-objective DTLZ, ZDT test set in Fig. 4, which are averaged over 21 independent runs. The legend is shown in the DTZL1 subfigure. From Fig. 4, we can see that when the number of function evaluations is less than 160, the proposed CI-EMO method converges relatively much faster for many problems, such as DTLZ1, DTLZ3, DTLZ5, DTLZ7, and ZDT1-4.
@@
Table 2: IGD+ STATISTIC RESULTS (MEAN AND STANDARD DEVIATION) FOR CI-EMO AND OTHER FIVE SAEAS ALGORITHMS ON MaF PROBLEMS
@@
4.3 Comparison with state-of-the-art SAEAs on Many-objective Problems
@@
To further research CI-EMO performance on expensive many-objective problems, we compare algorithms on 3-, 5-, and 10-objective MaF test problems. IGD+ Results are shown in Table 2. CI-EMO performs better than K-RVEA, KTA2, EMMOEA, DirHV-EGO, and R2/D-EGO on 29, 15, 15, 26, and 25 out of 33 problems, respectively, while being inferior to these algorithms on only 1, 5, 10, 4, and 4 problems, respectively. We find that CI-EMO is most efficient in solving MaF1-3 and MaF7-11. CI-EMO is weaker than EMMOEA on MaF4, MaF6, and MaF12-13 since EMMOEA has an advantage in solving concave PF problems, which has been described on (Qin et al., 2023). Table S-II shows the mean HV values and standard deviations obtained by the six algorithms. CI-EMO performs better than K-RVEA, KTA2, EMMOEA, DirHV-EGO, and R2/D-EGO on 27, 16, 11, 18, and 22 out of 33 problems, while it is inferior to them only on 0, 4, 9, 6, and 3 problems, respectively. Overall, CI-EMO has the best performance, which verifies the effectiveness of composite indicator-based sampling in high-dimensional objective space. Thus, experiments demonstrate CI-EMO not only has strong competitiveness in solving expensive multi-objective optimization problems but also works well on many-objective optimization problems.
4.4 Effectiveness of Composite Indicator
@@
To validate the effectiveness of the composite indicator, this paper designs two types of experiments: one to analyze the properties of each component in the composite indicator, and another to assess the impact of different components of the indicator by ablation experiments. First, CI-EMO is compared with four variants, where $rand$-EMO refers to the method of randomly selecting an individual from candidate population for real fitness evaluation, while the other three variants correspond to using one of the three performance indicators for candidate selection, respectively. “M” represents the number of objectives and "D" means the dimension of the problem. $\mathbf{I}_{1}$-EMO means only $\mathbf{I}_{1}$ is used in the composite indicator. The experimental results are presented in Table 3. As shown in the table, CI-EMO significantly outperforms $rand$-EMO on 13 test problems. Furthermore, the algorithm’s performance is inferior to CI-EMO when only a single indicator ($\mathbf{I}_{1}$, $\mathbf{I}_{2}$, or $\mathbf{I}_{3}$) is used. It is evident that using different indicators produces significantly varied results across different test problems, as shown in Fig. S-1. Compared with single sampling indicator-based variants, CI-EMO has better robustness. Among the three variants that use a single performance indicator, $\mathbf{I}_{1}$-EMO performs the best. It shows no significant difference compared to CI-EMO in 14 test cases and is significantly worse in five test problems. $\mathbf{I}_{2}$-EMO, performs significantly worse than CI-EMO in 9 test cases. It performs well on the ZDT test set but underperforms on the DTLZ test set. $\mathbf{I}_{3}$-EMO exhibits the poorest performance; except for DTLZ1 and DTLZ3, it is significantly worse than CI-EMO in most of the problems. As shown in Fig. S-1, if only $\mathbf{I}_{3}$ is used, solutions on PF will lack diversity, selection of solutions based on the convergence indicator is biased towards certain regions of the PF as results on DTLZ2 and ZDT3. In addition, the algorithm benefits from multiple iterations of surrogate model-assisted NSGA-III during the generation process of the candidate population. Individuals in the candidate population have exhibited convergence. Therefore, $\mathbf{I}_{1}$-EMO and $\mathbf{I}_{2}$-EMO have better performance than $\mathbf{I}_{3}$-EMO.
@@
On the other hand, ablation experiments have been conducted. We delete one metric from the composite indicator in each variant. Results have been shown in Table 4. CI-EMO-no-$\mathbf{I}_{1}$ means that $\mathbf{I}_{2}$ and $\mathbf{I}_{3}$ are used in the composite indicator, and $\mathbf{I}_{1}$ is deleted from the composite indicator. We can see that no matter which of the three metrics is deleted, the performance of the algorithm will drop significantly. Among them, deleting $\mathbf{I}_{3}$ has the least impact on the proposed algorithm, and the effect of CI-EMO-no-$\mathbf{I}_{3}$ is better on some problems, such as ZDT1-3.
@@
Table 3: IGD+ STATISTIC RESULTS (MEAN AND STANDARD DEVIATION) FOR CI-EMO AND OTHER FOUR VARIANTS ON BENCHMARK PROBLEMS
@@
Table 4: IGD+ STATISTIC RESULTS (MEAN AND STANDARD DEVIATION) FOR CI-EMO AND OTHER THREE VARIANTS ON BENCHMARK PROBLEMS
@@
Table 5: IGD+ STATISTIC RESULTS (MEAN AND STANDARD DEVIATION) FOR CI-EMO AND THE VARIANT WITHOUT NORMALIZATION
@@
4.5 Effectiveness of normalization
For integrating the three kinds of performance metrics of convergence, diversity, and distribution, the algorithm normalizes different indicators and then superimposes them together. To verify the impact of normalization on each indicator, we used a new variant CI-EMO-no-Norm as a control. In CI-EMO-no-Norm, $\mathbf{I}_{1}$, $\mathbf{I}_{2}$, and $\mathbf{I}_{3}$ are obtained without normalization, that is, they are equal to $\theta(\boldsymbol{x})$, $d_{c}\left(\boldsymbol{x}\right)$, and $-d_{z}\left(\boldsymbol{x}\right)$, respectively. The experimental result is shown in Table 5. CI-EMO significantly outperformed CI-EMO-no-Norm on 11 test questions, and there was no significant difference on the remaining 8 test questions. It shows that normalization can help combine different scale metrics and achieve the balance of convergence, diversity, and distribution.
Table 6: IGD+ STATISTIC RESULTS (MEAN AND STANDARD DEVIATION) FOR CI-EMO AND OTHER FOUR VARIANTS WITH DIFFERENT SAMPLING NUMBER
@@ 4.5 Effectiveness of normalization
4.6 Effect of random weights of indicators
A fixed weight setting may cause the algorithm’s sampling points to exhibit a single preference and lack diversity. To address this, random weights are assigned to the three sampling indicators during their integration. To analyze the impact of random weights, we compare CI-EMO with its variant, CI-EMO-SW, where all weights are fixed and set to 1. The results are presented in Table S-III. Experimental comparisons reveal that the overall statistical results of fixed weights and random weights are not significantly different. However, random weights yield more optimal results, as highlighted in Table S-III. Furthermore, as shown in Fig. 5, for the discontinuous Pareto front problem ZDT3, using random weights demonstrates better robustness in identifying each segment of the discontinuous Pareto front. Experiments show random weights enhance, improve the diversity of sampling, and prevent issues due to fixed weights.
@@ 4.6 Effect of random weights of indicators
4.7 Effect of sampling number setting each generation
In expensive optimization, the sampling number in each iteration is commonly an important parameter for algorithm performance. Some previous methods use multiple sampling strategies to select multiple candidate solutions from the candidate population at each iteration for real fitness evaluation, and each strategy is used for different purposes (Chugh et al., 2016; Song et al., 2021). In CI-EMO, we use a composite indicator to select only one candidate in each iteration, where the selected candidate balances the performance of convergence, diversity, and distribution.
To verify the impact of the sampling number of CI-EMO in this paper, we compared CI-EMO with four variants, where CI-EMO-q10 means the algorithm selects 10 candidate solutions for real fitness evaluation in each iteration. Similarly, CI-EMO-q5, CI-EMO-q3, and CI-EMO-q2 respectively indicate that 5, 3, and 2 candidate solutions are selected from the candidate population for real fitness value evaluation in each iteration. When the former candidate solution is selected based on the composite indicator, the candidate solution is deleted from the candidate population, and the method selects the next new candidate solution. As shown in Table 6, CI-EMO has the best performance. The performance of CI-EMO-q2 and CI-EMO-q3 are similar to CI-EMO. CI-EMO-q3 performs significantly worse than CI-EMO on only 3 test problems. As the number of samples continues to increase, the performance of the algorithm decreases significantly. This is because some unnecessary candidate solutions in the population are selected, making the use of the true fitness value evaluation times no longer efficient.
4.8 Computational Complexity Analysis